# Data-Driven Discovery of Digital Twins in Biomedical Research

### v1

@@ Abstract
Recent technological advances have expanded the availability of high-throughput biological datasets, enabling the reliable design of digital twins of biomedical systems or patients. Such computational tools represent key chemical reaction networks driving perturbation or drug response and can profoundly guide drug discovery and personalized therapeutics. Yet, their development still depends on laborious data integration by the human modeler, so that automated approaches are critically needed. The successes of data-driven system discovery in Physics, rooted in clean datasets and well-defined governing laws, have fueled interest in applying similar techniques in Biology, which present unique challenges. Here, we reviewed existing methodologies for automatically inferring digital twins from biological time series, which mostly involve symbolic or sparse regression. We evaluate algorithms according to eight biological and methodological challenges, associated to noisy/incomplete data, multiple conditions, prior knowledge integration, latent variables, high dimensionality, unobserved variable derivatives, candidate library design, and uncertainty quantification. Upon these criteria, sparse regression generally outperformed symbolic regression, particularly when using Bayesian frameworks. We further highlight the emerging role of deep learning and large language models, which enable innovative prior knowledge integration, though the reliability and consistency of such approaches need to be improved. While no single method addresses all challenges, we argue that progress in learning digital twins will come from hybrid and modular frameworks combining chemical reaction network-based mechanistic grounding, Bayesian uncertainty quantification, and the generative and knowledge integration capacities of deep learning. To support their development, we further propose a benchmarking framework to evaluate methods across all challenges.
@@ 1 Introduction
Recent technological advances have made increasingly available multi-type datasets documenting biological variables at multiple scales (subcellular, single cell, bulk, patient), over a dynamical time window, and under different experimental conditions. To handle the complexity of such data, several statistical and mathematical approaches have been proposed, from mechanism-based models of regulatory networks that enable straightforward validation procedures, to black-box statistical models (e.g., neural networks), which generally offer limited possibilities of exploring predictive features. Thus, medical research leans towards mechanistic models that represent the underlying biology and predict the molecular mechanisms driving observed phenotypes. Such digital twins of the patient or biological system constitute a powerful tool for the identification of disease-targetable vulnerabilities and personalized therapies. However, their design, i.e, the discovery of their governing equations, may be highly time-consuming for the human modeler as it requires integrating multi-type datasets with prior knowledge of various forms; this motivated the development of automatic model learning approaches.
@@
Model learning for Biology began with network inference, which aims to identify only gene interactions without capturing the system dynamics, a common example being Gene Regulatory Network (GRN) inference from transcriptomics  Marku and Pancaldi (2023). However, Biology may be better represented by quantitative models describing time-varying intracellular events using nonlinear ordinary differential equations (ODEs), where state variables are genes, proteins, or drug undergoing chemical reactions governed by specified laws and rate constants (Machado et al., 2011). Such a mathematical framework has been successfully applied to dynamically represent biological systems or drug pharmacokinetics-pharmacodynamics (PK-PD), thus populating the fields of systems biology and systems pharmacology (Kitano, 2002). Learning an ODE model implies inferring both i) the mathematical formulations of species interactions (e.g., law of mass action, Michaelis-Menten Keener and Sneyd (2008)) and ii) the magnitude of the reactions (e.g., rate or enzymatic constants). Traditionally, this process is carried out manually: the modeler reviews the literature, formulates equations, and fits parameters to data. However, data collection and integration may be time-consuming, and such an approach is only feasible when the reaction network is known, which may be the central question of the project. Moreover, manual model inference presents reproducibility issues, as different modelers may yield different models. Hence, there is a need for a systematic approach to mechanistic model learning.
In this review, we focus on methodologies inferring ODE-based digital twins from time-resolved biological datasets, which may document different cell types/patients and conditions (e.g., gene knockdown, drug treatment). We categorized methods as either symbolic or sparse regression algorithms (North et al., 2023) and recapitulated them in a ShinyApp  [^1] https://u1331systemspharmacology.shinyapps.io/model_learning_review/. Emphasis was put on methods amenable to discovering complex systems and on innovative approaches combining classical regression with artificial intelligence (AI) tools (Ghadami and Epureanu, 2022). Next, we identified four experimental and four methodological challenges currently associated with automatic learning of digital twins for Biology and Medicine, reported the performance of the most relevant methods and discussed the strategies to further manage them.
@@ 2.1 Problem Formulation
We are interested in inferring both the structure and the quantitative dynamics of an ODE-based model representing a biological system, capturing not only species interactions but also their precise mechanisms, and associated mathematical terms and parameters. Let us consider a biological system involving $m$ species, whose concentration vector $\bm{y}\in\mathbb{R}_{+}^{m}$ is experimentally observed on a discrete time grid provided as a dataset $\mathcal{D}=\{(t_{j},\bm{y}_{j})\}_{j=1}^{n}$. The dynamics of the state variables $\bm{x}(t)\in\mathbb{R}_{+}^{m}$ are modeled through the following system of ODEs parametrized by $\theta\in\mathbb{R}_{+}^{p}$ :
@@
where $\bm{\varepsilon}_{j}$ is the measurement error at observation time $t_{j}$. The key goal of model inference is to learn the vector field $\bm{f}:\mathbb{R}^{m}_{+},\mathbb{R}_{+}^{p}\to\mathbb{R}^{m}$ and parameters $\theta$ that best fit the provided datasets. Crucially, we seek to obtain a mechanistic estimator $\hat{\bm{f}}$ of $\bm{f}$, composed of biologically grounded terms, typically derived from the mass action law, Michaelis-Menten, or Hill kinetics.
@@ 2.2.1 Symbolic Regression
These methods typically rely on expression trees whose nodes are either a state variable, a constant or a basic operator, selected in a user-defined set (e.g., $\times,\div,\exp,\log$) and whose vertices indicate the link between those items leading to explicit mathematical formulas (Figure 1). The algorithms follow an iterative process, updating $\mathcal{P}_{i}^{l}=\left\{\hat{f}^{(i)}_{1},\hat{f}^{(i)}_{2},\dots,\hat{f}^{(i)}_{n_{i}}\right\}$ the current population of candidate expression trees for estimating the derivative of the species $x_{l}$ as follows. At iteration $i$, $\mathcal{P}_{i-1}^{l}$ undergoes perturbations, such as mutations and crossovers for Genetic Programming-based methods, to generate new expressions, which are then evaluated against the data, and only top performers $\mathcal{P}_{i}^{l}$ are used to seed the next generation (Algorithm 1). Candidate expressions are typically refined via evolutionary algorithms, Monte Carlo tree search, probabilistic grammars, or neural-symbolic models (La Cava et al., 2021).
@@
Output: Best symbolic model estimator $\hat{f}$ with adjusted parameters $\xi^{*}$
Step 1: Initialize the candidate models population $\mathcal{P}_{0}$ by:
@@
 for $k=1$ to $n_{i}$ do
      Fit free parameters of $\hat{f}^{(i)}_{k}\in\mathcal{P}_{i}$ using numerical optimization (e.g., least-squares), yielding the best-fit parameter set $\xi^{*(i)}_{k}=\arg\min_{\xi}\;L\!\left(\hat{f}^{(i)}_{k}(\cdot;\xi),\mathcal{D}\right).$
@@
 for $k=1$ to $n_{i}$ do
      Compute optional complexity score $C\left(\hat{f}^{(i)}_{k}(\cdot;\xi^{*(i)}_{k})\right)$ (e.g., Minimum Description Length or Pareto filtering)
@@
$J^{(i)}_{k}\;=\;L\!\left(\hat{f}^{(i)}_{k}(\cdot;\xi^{*(i)}_{k}),\mathcal{D}\right)\;+\;\lambda\,C\!\left(\hat{f}^{(i)}_{k}(\cdot;\xi^{*(i)}_{k})\right)$
    Retain the subset $\mathcal{S}_{i}\subset\mathcal{P}_{i}$ of top-performing candidates with the smallest $J^{(i)}_{k}$ for the next iteration
@@
Return final model: $\hat{f}(.;\xi^{*})\;=\;\arg\min_{k\in\{1,\dots,n_{N}\}}\left[L\left(\hat{f}_{k}^{(N)}(\cdot;\xi^{*(N)}_{k}),\mathcal{D}\right)+\lambda\,C\left(\hat{f}_{k}^{(N)}(\cdot;\xi^{*(N)}_{k})\right)\right$]
Algorithm 1 General Symbolic Regression Algorithm, encompassing most current approaches (evolutionary, neural, grammar-based, or constraint-guided), outlined for a 1-D problem. It presents the iterative generation and refinement of symbolic expression populations through candidate initialization, selection, generation, and optional post-processing.
@@
Figure 2 provides an overview of the symbolic regression methods we identified, which we grouped into six major families based on their core algorithmic strategies. First, genetic programming (GP) methods iteratively evolve candidate mathematical expressions using operators such as mutation and crossover (Figure 1), which are then selected over generations to minimize prediction error on data, often incorporating structural constraints or complexity penalties. The earliest contribution of such type is by Bongard and Lipson (2007) and interestingly combines (i) partitioning, i.e., modeling each variable independently; (ii) automated model probing, suggesting experimental perturbations (e.g., changes in initial conditions or parameters) to diversify the inferred models; and (iii) snipping, i.e., simplifying expressions by pruning redundant or low-impact components to enhance interpretability. This method was successfully tested on the 3-D lac operon regulatory network in E. coli with the best inferred model recovering 70% target terms. Next, neural-symbolic regression combines neural networks with symbolic modeling, either via analytical activation functions or symbolic decoders (i.e., neural modules that translate internal representations into symbolic expressions) that serve to generate candidate tree expressions at each iteration. More advanced strategies are also employed to constrain the symbolic search: reinforcement learning is used to efficiently explore the space of expressions, and neural networks can embed prior knowledge to guide the selection and refinement of candidate expressions. For example, ODEFormer (d’Ascoli et al., 2024) treats equation recovery as a symbolic translation task: it encodes time series trajectories and directly generates the corresponding symbolic expressions using a transformer decoder. It was evaluated on biologically relevant examples, including the SEIR epidemic model and the Lotka–Volterra system (4 and 2 variables, respectively), and showed strong performance and robustness even with noisy or limited data. Once pretrained, inference takes only a few seconds. Large Language Models (LLMs) further extend this idea by generating candidate equation structures via prompts. Recent applications have shown strong results on biological systems, including tumor growth and pharmacokinetics (Holt et al., 2024; Du et al., 2024; Shojaee et al., 2024). Grammar-based methods further rely on a formal specification of symbolic expression production rules as probabilities guiding the generation of new equation terms to be tested in the subsequent iteration. Brence et al. (2023) introduced grammar-guided sampling with unit-consistent constraints, enabling recovery of equations even with partial observability. Relevant to our current topic, Omejc et al. (2024) extended the method to ODE discovery from real-world data, which successfully recovered the dynamics of bacterial respiration and predator–prey systems, even with partial observability or low-frequency, noisy data. Next, Bayesian symbolic regression infers posterior distributions of expressions by combining structural priors and data likelihoods. For instance, Galagali and Marzouk (2019) employed reversible-jump MCMC to infer both network topology and reaction rate parameters, successfully identifying nonlinear mechanisms such as EGF–BRaf signaling. In a five-dimensional inference problem, their algorithm recovered the terms with a probability of about 0.75. Alternatively, mathematical constraint-based methods guide the model search by incorporating structural or domain-specific priors. Some methods leverage data properties or known scientific constraints (e.g., symmetries, conservation laws), others generate symbolic models from first principles (e.g., dimensional analysis), and a third group imposes structured search constraints based on equation syntax or consistency with trajectory behavior, reducing reliance on random exploration. Finally, of particular importance for the focus of this review, application-specific methods tailor symbolic regression to biological or chemical systems. Schmidt et al. (2011) refined ODE models of metabolic pathways by optimizing functional forms under biochemical constraints, such as stoichiometry and conservation laws. Reactmine (Martinelli et al., 2022) sequentially identifies likely reactions, such as synthesis, degradation, or enzyme catalysis, based on preponderant changes in species concentrations. The method inferred key known reactions in real-world datasets, one involving cell-cycle videomicroscopy, and another, circadian regulation based on gene expression and biomarker data (3 and 5 variables, respectively). On a synthetic MAPK phosphorylation cascade (7 variables), 6 out of 7 underlying reactions were recovered.
@@
Several benchmarks evaluate symbolic regression methods, starting from Orzechowski et al. (2018) who focused on low-dimensional algebraic expressions using noise-free data, and included only GP algorithms. Among those methods, EPLEX-1M (La Cava et al., 2016a) achieved the best generalization accuracy across 94 real-world datasets. Next, SRBench (La Cava et al., 2021) expanded this first benchmark by further incorporating neural methods and tests on dynamical systems, and DSR (Petersen et al., 2021) stood out as the best-performing method in terms of accuracy and expression simplicity on both dynamical and algebraic tasks. SRBench++ (de Franca et al., 2024) extended the approach by including controlled noise and unified metrics, yet focusing on algebraic equations. In this last comparative study, PySR (Cranmer, 2023) and Bingo (Randall et al., 2022) achieved the best overall harmonic trade-off between symbolic accuracy, robustness to noise, and expression simplicity, with uDSR Landajuela et al. (2022) showing perfect recovery on simpler tasks. Interestingly, PySR and DSR were also the best-performing methods at recovering chaotic systems, PySR being significantly faster and more robust on average (Gilpin (2021)). Next, ODEBench (d’Ascoli et al., 2024) presents the most structured benchmark for ODE modeling, testing 63 systems, with only 12 of them involving more than 3 variables. ODEFormer (d’Ascoli et al., 2024) consistently achieved the highest reconstruction and generalization accuracy across noise levels, while maintaining low inference time and expression complexity.
@@
Finally, Shojaee et al. (2025) introduced LLM-SRBench, a benchmark specifically designed to evaluate LLM-based symbolic regression methods. Instead of reusing equations, the authors generate synthetic one-dimensional equations by randomly assembling interpretable components from canonical biological mechanisms, like Hill functions, enriched with nonlinearities and structural diversity. This design ensures novelty and prevents trivial memorization by pretrained models. The benchmark includes 128 equations from various scientific domains, each paired with simulated time-series data, and evaluates symbolic coherence, numerical accuracy, and extrapolation. LLM-SR methods outperform traditional baselines for such metrics  (Shojaee et al., 2024), but the best-performing inferred model achieved only 31.5% symbolic accuracy.
@@
Overall, symbolic regression offers flexible formulations enabling the discovery of models faithfully representing the data, with the need for minimal prior knowledge, but at the cost of high computational time. Existing benchmarks have enabled meaningful methods evaluation by highlighting important trade-offs between accuracy, interpretability, and generalization. Still, most benchmarks rely on synthetic, low-dimensional, and fully observed systems, often with known ground truths inside the search space, which can overstate performance.

### v2

@@ Abstract
Recent technological advances have expanded the availability of high-throughput biological datasets, enabling the reliable design of digital twins of biomedical systems or patients. Such computational tools represent key chemical reaction networks driving perturbation or drug response and can profoundly guide drug discovery and personalized therapeutics. Yet, their development still depends on laborious data integration by the human modeler, so that automated approaches are critically needed. The successes of data-driven system discovery in Physics, rooted in clean datasets and well-defined governing laws, have fueled interest in applying similar techniques in Biology, which presents unique challenges. Here, we reviewed existing methodologies for automatically inferring digital twins from biological time series, which mostly involved symbolic or sparse regression. We evaluated algorithms according to eight biological and methodological challenges, associated to integrating noisy/incomplete data, multiple conditions, prior knowledge, latent variables, or dealing with high dimensionality, unobserved variable derivatives, candidate library design, and uncertainty quantification. Upon these criteria, sparse regression generally outperformed symbolic regression, particularly when using Bayesian frameworks. Next, deep learning and large language models, further emerge as innovative tools to integrate prior knowledge, although their reliability and consistency need to be improved. While no single method addresses all challenges, we argue that progress in learning digital twins will come from hybrid and modular frameworks combining chemical reaction network-based mechanistic grounding, Bayesian uncertainty quantification, and the generative and knowledge integration capacities of deep learning. To support their development, we further propose a benchmarking framework to evaluate methods across all challenges.
@@ 1 Introduction
Recent technological advances have made increasingly available multi-type datasets documenting biological variables at multiple scales (subcellular, single cell, bulk, patient), over a dynamical time window, and under different experimental conditions. To handle the complexity of such data, several statistical and mathematical approaches have been proposed, from mechanism-based models of regulatory networks that enable straightforward validation procedures, to black-box statistical models (e.g., neural networks), which generally offer limited possibilities of exploring predictive features. Thus, medical research leans towards mechanistic models that represent the underlying biology and predict the molecular mechanisms driving observed phenotypes. Such digital twins of the patients or biological systems constitute a powerful tool for the identification of disease-targetable vulnerabilities and personalized therapies. However, their design, i.e, the discovery of their governing equations, may be highly time-consuming for the human modeler as it requires integrating multi-type datasets with prior knowledge of various forms; this motivated the development of automatic model learning approaches.
@@
Model learning for Biology began with intracellular network discovery, which aims to only identify gene interactions without capturing the system dynamics, a common example being Gene Regulatory Network (GRN) inference from transcriptomics  Marku and Pancaldi (2023). However, Biology may be better represented by quantitative models describing time-varying intracellular events using nonlinear ordinary differential equations (ODEs), where state variables are genes, proteins, or drug species undergoing chemical reactions governed by specified laws and rate constants (Machado et al., 2011). Such a mathematical framework has been successfully applied to dynamically represent biological systems or drug pharmacokinetics-pharmacodynamics (PK-PD), thus populating the fields of systems biology and systems pharmacology (Kitano, 2002). Learning an ODE model implies inferring both i) the mathematical formulations of species interactions (e.g., law of mass action, Michaelis-Menten Keener and Sneyd (2008)) and ii) the magnitude of the reactions (e.g., rate or enzymatic constants). Traditionally, this process is carried out manually: the modeler reviews the literature, formulates equations, and fits parameters to data. However, data collection and integration may be time-consuming, and such an approach is only feasible when the reaction network is known, which may be the central question of the project. Moreover, manual model inference presents reproducibility issues, as different modelers may yield different models. Hence, there is a need for a systematic approach to mechanistic model learning.
In this review, we focus on methodologies inferring ODE-based digital twins from time-resolved biological datasets, which may document different cell types/patients and conditions (e.g., gene knockdown, drug treatment). We categorized methods as either symbolic or sparse regression algorithms (North et al., 2023) and recapitulated them in a ShinyApp  [^1] https://u1331systemspharmacology.shinyapps.io/model_learning_review/. Emphasis was put on methods amenable to discovering complex systems and on innovative approaches combining classical regression with artificial intelligence (AI) tools (Ghadami and Epureanu, 2022). We identified four experimental and four methodological challenges currently associated with automatic learning of digital twins for Biology and Medicine, reported the performance of the most relevant methods and discussed the strategies to further manage them.
@@ 2.1 Problem Formulation
We are interested in inferring both the structure and the quantitative dynamics of an ODE-based model representing a biological system, capturing not only species interactions but also their precise mechanisms, and associated mathematical terms and parameters. Let us consider a biological system involving $m$ species, whose concentration vector $\bm{y}\in\mathbb{R}_{+}^{m}$ is experimentally observed on a discrete time grid provided as a dataset $\mathcal{D}=\{(t_{j},\bm{y}_{j})\}_{j=1}^{n}$. The dynamics of the state variables $\bm{x}(t)\in\mathbb{R}_{+}^{m}$ are modeled through the following system of ODEs parametrized by $\theta\in\mathbb{R}_{+}^{s}$ :
@@
where $\bm{\varepsilon}_{j}$ is the measurement error at observation time $t_{j}$. The key goal of model inference is to learn the vector field $\bm{f}:\mathbb{R}^{m}_{+},\mathbb{R}_{+}^{s}\to\mathbb{R}^{m}$ and parameters $\theta$ that best fit the provided datasets. Crucially, we seek to obtain a mechanistic estimator $\hat{\bm{f}}$ of $\bm{f}$, composed of biologically grounded terms, typically derived from the mass action law, Michaelis-Menten, or Hill kinetics.
@@ 2.2.1 Symbolic Regression
These methods typically rely on expression trees whose nodes are either a state variable, a constant or a basic operator, selected in a user-defined set (e.g., $\times,\div,\exp,\log$) and whose vertices indicate the link between those items leading to explicit mathematical formulas (Figure 1). The algorithms iteratively optimize the populations of candidate expression trees for each state variable independently. At iteration $i$, the $n_{i}$ candidate functions for estimating $x_{l}$ derivative, $\mathcal{P}_{i}^{l}=\left\{\hat{f}^{(i)}_{1},\hat{f}^{(i)}_{2},\dots,\hat{f}^{(i)}_{n_{i}}\right\}$, undergoes perturbations, such as mutations and crossovers for Genetic Programming-based methods, to generate new expressions, which are then evaluated against the data, and only top performers are retained to form $\mathcal{P}_{i+1}^{l}$ (Algorithm 1). Candidate expressions are typically refined via evolutionary algorithms, Monte Carlo tree search, probabilistic grammars, or neural-symbolic models (La Cava et al., 2021).
@@
Output: Best symbolic model estimator $\hat{f}$ with adjusted parameters $\theta^{*}$
Step 1: Initialize the candidate models population $\mathcal{P}_{0}$:
@@
 for $k=1$ to $Card(\mathcal{P}_{i})$ do
      Fit free parameters of $\hat{f}^{(i)}_{k}\in\mathcal{P}_{i}$ using numerical optimization (e.g., least-squares), yielding the best-fit parameter set $\theta^{*(i)}_{k}=\arg\min_{\theta^{(i)}_{k}}\;L\!\left(\hat{f}^{(i)}_{k}(\cdot;\theta^{(i)}_{k}),\mathcal{D}\right).$
@@
 for $k=1$ to $Card(\mathcal{P}_{i})$ do
      Compute optional complexity score $C\left(\hat{f}^{(i)}_{k}(\cdot;\theta^{*(i)}_{k})\right)$ (e.g., Minimum Description Length or Pareto filtering)
@@
$J^{(i)}_{k}\;=\;L\!\left(\hat{f}^{(i)}_{k}(\cdot;\theta^{*(i)}_{k}),\mathcal{D}\right)\;+\;\lambda\,C\!\left(\hat{f}^{(i)}_{k}(\cdot;\theta^{*(i)}_{k})\right)$
    Retain the subset $\mathcal{S}_{i}\subset\mathcal{P}_{i}$ of top-performing candidates with the smallest $J^{(i)}_{k}$
@@
Return final model: $\hat{f}(.;\theta^{*})\;=\;\arg\min_{k\in\{1,\dots,n_{N}\}}\left[L\left(\hat{f}_{k}^{(N)}(\cdot;\theta^{*(N)}_{k}),\mathcal{D}\right)+\lambda\,C\left(\hat{f}_{k}^{(N)}(\cdot;\theta^{*(N)}_{k})\right)\right$]
Algorithm 1 General Symbolic Regression Algorithm, encompassing most current approaches (evolutionary, neural, grammar-based, or constraint-guided), outlined for a 1-D problem. It presents the iterative process of symbolic regression through candidate populations initialization, selection, generation, and optional post-processing.
@@
Figure 2 provides an overview of the symbolic regression methods we identified, which we grouped into six major families based on their core algorithmic strategies. First, genetic programming (GP) methods iteratively evolve candidate mathematical expressions using operators such as mutation and crossover (Figure 1), which are then selected over generations to minimize the prediction error on data, often incorporating structural constraints or complexity penalties. The earliest contribution of such type is by Bongard and Lipson (2007) and interestingly combines (i) partitioning, i.e., modeling each variable independently; (ii) automated model probing, suggesting experimental perturbations (e.g., changes in initial conditions or parameters) to diversify the inferred models; and (iii) snipping, i.e., simplifying expressions by pruning redundant or low-impact components to enhance interpretability. This method was successfully tested on the 3-D lac operon regulatory network in E. coli with the best inferred model recovering 70% of the target terms. Next, neural-symbolic regression combines neural networks with symbolic modeling, via either analytical activation functions or symbolic decoders (i.e., neural modules that translate internal representations into symbolic expressions) that serve to generate candidate tree expressions at each iteration. More advanced strategies are also employed to constrain the symbolic search: reinforcement learning is used to efficiently explore the space of expressions, and neural networks can embed prior knowledge to guide the selection and refinement of candidate expressions. For example, ODEFormer (d’Ascoli et al., 2024) treats equation recovery as a symbolic translation task: it encodes time series trajectories and directly generates the corresponding symbolic expressions using a transformer decoder. It was evaluated on biologically relevant examples, including the SEIR epidemic model and the Lotka–Volterra system (4 and 2 variables, respectively), and showed strong performance and robustness even with noisy or limited data. Once the framework is pretrained, inference of such low-dimensional models takes only a few seconds. Large Language Models (LLMs) further extend this idea by generating candidate equation structures via prompts. Recent applications have shown strong results on biological systems, including tumor growth and pharmacokinetics (Holt et al., 2024; Du et al., 2024; Shojaee et al., 2024). Alternatively, grammar-based methods rely on a formal specification of symbolic expression production rules as probabilities guiding the generation of new equation terms to be tested in the subsequent iteration. Brence et al. (2023) introduced grammar-guided sampling with unit-consistent constraints, enabling recovery of equations even with partial observability. Relevant to our current topic, Omejc et al. (2024) extended the method to ODE discovery from real-world data, and successfully recovered the dynamics of bacterial respiration and predator–prey systems, even with partial observability or low-frequency and noisy data. Next, Bayesian symbolic regression infers posterior distributions of expressions by combining structural priors and data likelihoods. For instance, Galagali and Marzouk (2019) employed reversible-jump MCMC to infer both network topology and reaction rate parameters, and successfully identified nonlinear mechanisms of EGF–BRaf signaling. In a five-dimensional inference problem, their algorithm recovered the terms with a probability of about 0.75. Alternatively, mathematical constraint-based methods guide the model search by incorporating structural or domain-specific priors. Some methods leverage data properties or known scientific constraints (e.g., symmetries, conservation laws), others generate symbolic models from first principles (e.g., dimensional analysis), and a third group imposes structured search constraints based on equation syntax or consistency with trajectory behavior, reducing reliance on random exploration. Finally, of particular importance for this review, application-specific methods tailor symbolic regression to biological or chemical systems. Schmidt et al. (2011) refined ODE models of metabolic pathways by optimizing functional forms under biochemical constraints, such as stoichiometry and conservation laws. Reactmine (Martinelli et al., 2022) sequentially identifies likely reactions, such as synthesis, degradation, or enzyme catalysis, based on preponderant changes in species concentrations. The method inferred key known reactions in real-world datasets, one involving cell-cycle videomicroscopy, and another, circadian regulation based on gene expression and biomarker data (3 and 5 variables, respectively). On a synthetic MAPK phosphorylation cascade (7 variables), 6 out of 7 underlying reactions were recovered.
@@
Several benchmarks evaluate symbolic regression methods, starting from Orzechowski et al. (2018) who focused on low-dimensional algebraic expressions using noise-free data, and included only GP algorithms. Among those methods, EPLEX-1M (La Cava et al., 2016a) achieved the best generalization accuracy across 94 real-world datasets. Next, SRBench (La Cava et al., 2021) expanded this first benchmark by incorporating neural methods and tests on dynamical systems, and revealed DSR (Petersen et al., 2021) as the best-performing method in terms of accuracy and expression simplicity on both dynamical and algebraic tasks. SRBench++ (de Franca et al., 2024) extended the approach by including controlled noise and unified metrics, yet focusing on algebraic equations. In this last comparative study, PySR (Cranmer, 2023) and Bingo (Randall et al., 2022) achieved the best overall harmonic trade-off between symbolic accuracy, robustness to noise, and expression simplicity, with uDSR Landajuela et al. (2022) showing perfect recovery on simpler tasks. Interestingly, PySR and DSR were also the best-performing methods at recovering chaotic systems, PySR being significantly faster and more robust on average  (Gilpin, 2021). Next, ODEBench (d’Ascoli et al., 2024) presents the most structured benchmark for ODE modeling, testing 63 systems, yet with only 12 of them involving more than 3 variables. ODEFormer (d’Ascoli et al., 2024) consistently achieved the highest reconstruction and generalization accuracy across noise levels, while maintaining low inference time and expression complexity.
@@
Finally, Shojaee et al. (2025) introduced LLM-SRBench, a benchmark specifically designed to evaluate LLM-based symbolic regression methods. Instead of reusing equations, the authors generate synthetic one-dimensional equations by randomly assembling interpretable components from canonical biological mechanisms, like Hill functions, enriched with nonlinearities and structural diversity. This design ensures novelty and prevents trivial memorization by pretrained models. The benchmark includes 128 equations from various scientific domains, each paired with simulated time-series data, and evaluates symbolic coherence, numerical accuracy, and extrapolation. LLM-SR methods outperform traditional baselines for such metrics  (Shojaee et al., 2024), but the best-performing inferred model achieved a symbolic accuracy of only 31.5% .
@@
Overall, symbolic regression offers flexible formulations enabling the discovery of models faithfully representing the data, with the need for minimal prior knowledge, but currently tested only on models with a small number of variables. Existing benchmarks have enabled meaningful methods evaluation by highlighting important trade-offs between accuracy, interpretability, and generalization. Still, most benchmarks rely on synthetic, low-dimensional, and fully observed systems, often with known ground truths inside the search space, which can overstate performance.